A chemist needs a 130 millimeter of a 57% soulution but has only 33% and 85% soulutions available. Find how many millimeters of each that should be mixed to get the desired soulution.
Answer by Alan3354(69443)

A chemist needs a 130 millimeter of a 57% soulution but has only 33% and 85% soulutions available. Find how many millimeters of each that should be mixed to get the desired soulution.
---------------
Do it like this one:
if a 14% acid solution is mixed with a 19% acid solution to produce 45 gallons of a 15% acid solution how much (in gallons) of each solution should be used?
--------------
f = 14%
n = 19%
----
f + n = 45 (total gallons)
14f + 19n = 15*45 (total acid)
-----
n = 45 - f
Sub for n
14f + 19*(45-f) = 675
14f + 855 - 19f = 675
-5f = -180
f = 36 gallons of 14%
n = 9 gallons of 19%